OpenStudy (anonymous):

Am I able to evaluate sin(cos^1(7/11)) without a calculator? If yes, how?

OpenStudy (anonymous):

Is that supposed to be cosine inverse?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[\text{Sin}\left[\text{ArcCos}\left[\frac{7}{11}\right]\right]\]

OpenStudy (anonymous):

You want to draw the right triangle associated with that. Im trying to think of a better way to explain it >.<

OpenStudy (anonymous):

\[\cos^{-1}(\frac{7}{11}) \Rightarrow \cos(\theta) = \frac{7}{11}\] and in a right triangle: \[\cos(\theta) = \frac{adj}{hyp}\] You want to draw that right triangle.

OpenStudy (anonymous):

\[cos(\theta) = \frac{7}{11} = \frac{adj}{hyp}\] \[adj^2 + opp^2 = hyp^2\]

OpenStudy (anonymous):

\[sin(\theta) = \frac{opp}{hyp}\]

OpenStudy (anonymous):

so find opp using the Pythagorean triangle and then put it over the hypotenuse.

OpenStudy (anonymous):

6(rad2)/11